# Vitali Telezkia and Stefan Klumppa

### v1

@@ 3.1 No external field
As a reference, we first simulated active dipolar particles without an external magnetic field ($B=0$). Here, we briefly present our results, in particular how the observed states depend on the activity $v_{0}$ and the pairwise magnetic interactions via the magnetic moment $\mu$. This system has previously been analyzed in detail by Liao et al. 14 and the following results are in agreement with their results. In simulations for different $v_{0}$ and $\mu$ and different densities $\Phi$, we observe different patterns. Representative snapshots are shown in figure 1. These snapshots reflect combinations of three different types of ordering: alignment of the magnetic moments, clustering of the particles, and chain formation. To quantify the order in these patterns, we make use of combinations of the three previously introduced global order parameters, the fraction of particles in the largest cluster $\left\langle\phi_{c}\right\rangle$, the global polarization $\left\langle\phi_{e}\right\rangle$, and the degree of polymerization $\left\langle\phi_{p}\right\rangle$, (see Methods). All three parameters are dimensionless quantities with values between zero and one. For a systematic classification, we consider the system as ordered if the corresponding order parameter exceed 0.5. This choice defines three criteria for order and combinations of the three criteria thus define eight possible ordered or disordered states that are summarized in table 1.
@@
Fig.  1: Snapshots showing examples for states introduced in table 1 for systems at low density ($\Phi=0.13$). Snapshots were taken at time $t=20$. The following states are shown: a. disordered gas ($\mu=0$, $v_{0}=23$), b. oriented gas ($\mu=2.6$, $v_{0}=500$), c. gas of chains ($\mu=1.6$, $v_{0}=23$), d. oriented chains ($\mu=1.6$, $v_{0}=500$), e. network of chains ($\mu=2.6$, $v_{0}=23$), f. bands ($\mu=4$, $v_{0}=100$). The color wheel (top left) indicates the orientation of the particles.
@@
Fig.  2: Diagram of states in the absence of an external field for three densities $\Phi=$ 0.13, 0.23, 0.57 (left to right). The states result from order parameter criteria combinations according to table 1. The red box in the right diagram ($\Phi=0.57$) highlights cluster formation that corresponds to MIPS for apolar ($\mu=0$) active particles.
@@
At the low and intermediate densities we simulated, we can distinguish six of these states that are shown in the snapshots in figure 1: the disordered gas state with no order, the oriented gas with only orientational order, the gas of chains with only chain formation, oriented chains (chain formation and orientational order), a network of chains based on the combination of chain formation and clustering, and bands in the case where all three types of order are combined. At high system density, well above the critical density for motility induced phase separation (MIPS, $\Phi_{\textnormal{MIPS}}\approx 0.28$ 35), we also see cluster formation in the apolar limit ($\mu=0$) that corresponds to MIPS, as expected 14. The eighth possible state, oriented clusters, characterized by clustering and orientational order, but no chaining is not seen in the absence of an external field.
Based on the systematic classification, we summarized our results in the diagrams of states shown in figure 2. For systems with densities below the critical density of MIPS ($\Phi=0.13$ and $\Phi=0.23$, figure 2 left and center) and weak magnetic interactions ($\mu$ < 1) no classification criterion is fulfilled and the system is in a disordered gas state (gray circle in figure 2, snapshot in figure 1 a). Order (chain formation and/or global polarization) emerges for stronger magnetic interactions ($\mu\geq 1.0$). At this point, contributions from dipolar interactions equal the energy contributed by thermal noise ($\mu^{2}=T=1$) and induce chain formation, consistent with previous observations in small systems 15. Different ordered states can be realized by varying the activity $v_{0}$ for fixed magnetic moment. For example, in the case of intermediate magnetic interactions ($\mu=1.4$) systems with low levels of activity tend to form a gas of chains (pink triangle in figure 2, snapshot in figure 1 c). Increasing the activity adds global polarization to the chaining and, thus, results in oriented chains (green pentagon in figure 2, snapshot in figure 1 d) form. For even higher levels of activity, the formation of chains is suppressed while global polarization persists, so that an oriented gas is observed (purple plus in figure 2, snapshot in figure 1 b).
When the system density exceeds the critical density of MIPS ($\Phi=0.57$, figure 2 right), apolar ($\mu=0$) and weakly magnetic particles ($\mu$ < 1) also show ordered states like a network of chains (yellow square in figure 2, snapshot in figure 1 e) or a gas of chains (pink triangle). In case of apolar particles, increasing the activity results in cluster formation (brown triangle) which has previously been identified as MIPS 14 (and was confirmed here by a bimodal distributions of the local density). At high system density, network of chains (yellow square) and bands (red cross in figure 2, snapshot in figure 1 f) dominate the diagram of states.
3.2 Influence of an external field
@@
Fig.  3: Collection of snapshots of the system at low density ($\Phi=0.13$) for intermediate self-propulsion speed $v_{0}=23$ for different magnetic field strengths $B$ (rows) and values for magnetic interaction strengths $\mu$ (columns). External magnetic field points from left to right. All snapshots have been taken at time $t=20$.
Figure 3 shows typical system configurations for three different values of the magnetic moments $\mu=$ 1.0, 1.4, 2.6 (columns from left to right) at these three field strengths ($B=$ 0.1, 1.0, 14 in the top, middle, bottom row, respectively) for particles with moderate self-propulsion speed ($v_{0}=23$). The orientation of the particles is indicated by their color according to the color wheel. The magnetic field points from left to right, so alignment with the field can be seen by the cyan color of the particles. Systems with a weak external magnetic field (top row) display similar configurations to what was observed in systems without external magnetic field (compare with figure 2). In systems with intermediate external magnetic fields (middle row), particles align with the direction of the external field (as seen by their cyan color). In addition, particles assemble into chains for strong external magnetic fields ($B=14$). These chains often have defects such as branches or small loops. Here, the average chain length is seen to increase with the magnetic moment of the particles, spanning the whole system size for strong magnetic interactions and strong external field strengths (for $\mu=2.6,\,B=14$, bottom right). These columnar ordered clusters are qualitatively similar to structures formed by non-equilibrium ferrofluids with external magnetic fields 23. In contrast to systems without an external magnetic field, these columnar clusters mostly stay evenly separated and do not combine into connected networks of chains or into broad bands (at the low density simulated here, broad bands are seen for higher density, as discussed below). In fact, we observed in systems with strong magnetic interactions ($\mu=2.6$) that an increase of the external magnetic field strength causes broad bands to separate into thinner bands with fewer lanes or into individual chains of dipolar particles (compare the cases for $\mu=2.6$, $B=1.0$ and $\mu=2.6$, $B=14$ in figure 3).
We analyzed the numerical simulations quantitatively by calculating the three global order parameters introduced above. They are plotted in figure 4 for different $\mu$ and $v_{0}$ and for strong and weak magnetic field. Clustering (with $\left\langle\phi_{c}\right\rangle>0.5$) is largely absent at the low density used here and is only observed for the highest magnetic moment and a weak magnetic field (corresponding to the bands or networks of chains seen in the absence of a magnetic field). The strong magnetic field suppresses clustering also under these conditions (red lines).
@@
Fig.  4: Three order parameters: fraction of largest cluster (top), polarization (center) and polymerization (bottom) against self-propulsion speed $v_{0}$ for four magnetic moments $\mu$ (colors). Solid color and solid lines show data for strong external magnetic field $B=14$, transparent colors and dashed lines show data with weak magnetic field strength $B=0.1$. System density was set to $\Phi=0.13$. Cut-off for cluster formation set at $\epsilon=1.28$.
@@
Fig.  5: Diagrams of states for three different external magnetic field strengths $B=$ 0.1, 1.0, 14 (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at low density ($\Phi=0.13$).
@@
Using again the criteria based on the three order parameters, we classified the observed states of the system and summarized the results in a diagram of states, shown in figure 5. The figure shows the diagrams of states for the three different magnetic field strengths $B=$ 0.1 1.0 14 (left to right) and for the density $\Phi=0.13$. Based on these diagrams we can make three key observations: First, disordered gas states (gray circles) are strongly suppressed with increasing magnetic field strength. In addition, the onset of the oriented gas state (purple plus) shifts to lower values of $\mu$. Second, intermediate as well as strong magnetic fields (center and right diagram) suppress the formation of networks of chains (yellow square) and instead promote the formation of oriented states, specifically the oriented gas (purple plus), oriented chains (green pentagon) and bands (red cross, for intermediate field strength). Third, and most importantly, strong external magnetic fields suppress the formation of bands. For strong external magnetic fields (right diagram), two states dominate the diagram, the oriented gas (purple plus) and oriented chains (green pentagon). The line separating these two states resembles the corresponding line separating the gas and chain state in small confined systems of dipolar active particles investigated in previous work 15.
@@
For strong magnetic fields, band patterns were largely absent (right panel in figure 5). We only found bands for two parameter combinations ($\mu=2.6$, $v_{0}=200$ and $\mu$=4, $v_{0}=500$, indicated by the red crosses in figure 5). In these cases we observed the following noteworthy behavior, shown by the snapshots in figure 6. Dipolar active particles that have self-assembled into long chains transiently show oscillations. The amplitude of these oscillations was seen to grow towards their tail of a chain. We suspect that such oscillations might be caused by a buckling instability: In the chains, particles with strong dipolar interactions are packed rather closely. The combination of high swimming speeds and translational noise can then cause a temporary compression and stress along the chain that is resolved by buckling of the chain. When the dipolar interactions were reduced, long chains were seen to break into smaller segments. These breaks typically occurred near the tail of the chain, where the oscillations reach their maximum amplitude. This observation suggests that fragmentation induced by buckling may provide a mechanism for the coexistence of chains of different lengths. We note that the breaking of chains resembles the fragmentation of passive chains observed in magnetorheological suspensions driven by a rotating field 36, 37. In our simulations these oscillating chains are sees as a transient state towards band formation. At some point, individual oscillating chains meet. Then, alignment with the neighboring particles suppresses oscillations and a band starts to form. Again, similar lateral aggregation of individual chains into bands and columnar lanes has been observed in magnetorheological fluids 22, 23, 24, 38.
@@
Fig.  6: Series of snapshots showing oscillating chains for dipolar active particles with strong magnetic interactions $\mu=4$ and high self-propulsion speed $v_{0}=500$. Snapshots taken at times 26.46, 26.56, 26.66 (left to right), color wheel indicates orientations. External magnetic field points from left to right, magnetic field strength was set to $B=14$.
3.3 Effect of density
@@
To investigate the effect of the density of the system on the patterns formed by the active dipolar particles in a magnetic field, we simulated systems with two higher densities ($\Phi=0.23$ and $\Phi=0.57$) and repeated the calculation of the order parameters. The resulting state diagrams are shown in figure 7 (for $\Phi=0.23$) and in figure 8 (for $\Phi=0.57$).
@@
In systems with the intermediate density $\Phi=0.23$, we observed that intermediate external magnetic field strengths of $B=1$ do not suffice to suppress formation of complex networks (yellow squares, compare the middle panel in (figure 7 with that for low density in figure 5). Because of the increase in density and therefore a reduced mean particle distance, short chains may connect and form a complex network.
@@
Furthermore, the diagram of states indicates that strong external magnetic fields diminish the effect activity has in systems at intermediate density (figure 7, right). Here, the line separating the oriented gas state (purple plus) and the oriented chains state (green pentagon) is independent of $v_{0}$. This observation is markedly different from the previous observations for systems without external magnetic fields and to systems at low density with external magnetic fields, where activity had a strong influence on the collective behavior and structure formation.
@@
Fig.  7: Diagrams of states for three different external magnetic field strengths $B=$ 0.1, 1.0, 14 (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at intermediate density ($\Phi=0.23$). Cut-off for cluster formation set at $\epsilon=1.21$.
@@
For systems of high density (our case with $\Phi=0.57$), three phenomena compete with one another: polar order, chain formation and MIPS, as pointed out by Liao et al. 14. The addition of an external magnetic field introduces global polar order and suppresses formation of complex networks of chains. Instead, we observed that strong external magnetic field strengths of $B=14$ clearly organize the system with high magnetic moments into columnar structures (see figure 3 $\mu=2.6,B=14$). Because of the increase in density, these columnar structures are made up of multiple lanes in off-register alignment and not of single lanes, as seen in systems at low density.
@@
In the diagram of states, we observed that the increased density causes the previously suppressed band states (red crosses) to reappear. In fact, bands dominate the corresponding diagram of states for strong external magnetic fields and are independent of activity (figure 8, right). Off-register band formation has been previously observed in passive dipolar systems with constant external magnetic fields 39, 40, 41. In systems at high density, a clear discrimination between bands and columnar states can no longer be made. A limiting factor might be the finite system size, as observed in other systems of active particles 42.
@@
Fig.  8: Diagrams of states for three different external magnetic field strengths $B=$ 0.1, 1.0, 14 (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at high density ($\Phi=0.57$). Cut-off for cluster formation set at $\epsilon=1.05$.
3.4 Columnar clusters

### v2

@@ III.1 No external field
As a reference, we first simulated active dipolar particles without an external magnetic field ($B=0$). Here, we briefly present our results, in particular how the observed states depend on the activity $v_{0}$ and the pairwise magnetic interactions via the magnetic moment $\mu$. This system has previously been analyzed in detail by Liao et al. [14] and the following results are in agreement with their results. In simulations for different $v_{0}$ and $\mu$ and different densities $\Phi$, we observe different patterns. Representative snapshots are shown in fig. 2. These snapshots reflect combinations of three different types of ordering: alignment of the magnetic moments, clustering of the particles, and chain formation. To quantify the order in these patterns, we make use of combinations of the three previously introduced global order parameters, the fraction of particles in the largest cluster $\left\langle\phi_{c}\right\rangle$, the global polarization $\left\langle\phi_{e}\right\rangle$, and the degree of polymerization $\left\langle\phi_{p}\right\rangle$ (see Methods). All three parameters are dimensionless quantities with values between zero and one. For a systematic classification, we consider the system as ordered if the corresponding order parameter exceed 0.5. This choice defines three criteria for order and combinations of these three criteria thus define eight possible states, as summarized in table 1.
@@
Figure 2: Snapshots showing examples for states introduced in table 1 for systems at low density ($\Phi=0.13$), without an external field. Snapshots were taken at time $t=20$. The following states are shown: a. disordered gas ($\mu=0$, $v_{0}=23$), b. oriented gas ($\mu=2.6$, $v_{0}=500$), c. gas of chains ($\mu=1.6$, $v_{0}=23$), d. oriented chains ($\mu=1.6$, $v_{0}=500$), e. network of chains ($\mu=2.6$, $v_{0}=23$), f. bands ($\mu=4$, $v_{0}=100$). The color wheel (top left) indicates the orientation of the particles.
@@
Figure 3: Diagram of states in the absence of an external field for three densities $\Phi=$0.13 , 0.23 and 0.570.130.230.570.130.230.57start_ARG 0.13 end_ARG , start_ARG 0.23 end_ARG and start_ARG 0.57 end_ARG (left to right). The states result from order parameter criteria combinations according to table 1. The red box in the right diagram ($\Phi=0.57$) highlights cluster formation that corresponds to MIPS for apolar ($\mu=0$) active particles.
@@
At the low and intermediate densities we simulated, we can distinguish six of these states that are shown in the snapshots in fig. 2: the disordered gas state with no order, the oriented gas with only orientational order, the gas of chains with only chain formation, oriented chains (chain formation and orientational order), a network of chains based on the combination of chain formation and clustering, and bands in the case where all three types of order are combined. At high system density, well above the critical density for motility induced phase separation (MIPS, $\Phi_{\textnormal{MIPS}}\approx 0.28$ [35]), we also see cluster formation in the apolar limit ($\mu=0$) that corresponds to MIPS, as expected [14]. The eighth possible state, oriented clusters, characterized by clustering and orientational order, but no chaining is not seen in the absence of an external field.
Based on the systematic classification, we summarized our results in the diagrams of states shown in fig. 3. For systems with densities below the critical density of MIPS ($\Phi=0.13$ and $\Phi=0.23$, fig. 3 left and center) and weak magnetic interactions ($\mu$ 1) no classification criterion is fulfilled and the system is in a disordered gas state (gray circle in fig. 3, snapshot in fig. 2 a). Order (chain formation and/or global polarization) emerges for stronger magnetic interactions ($\mu\geq 1.0$). At this point, contributions from dipolar interactions equal the energy contributed by thermal noise ($\mu^{2}=T=1$) and induce chain formation, consistent with previous observations in small systems [15]. Different ordered states can be realized by varying the activity $v_{0}$ for fixed magnetic moment. For example, in the case of intermediate magnetic interactions ($\mu=1.4$) systems with low levels of activity tend to form a gas of chains (pink triangle in fig. 3, snapshot in fig. 2 c). Increasing the activity adds global polarization to the chaining and, thus, results in oriented chains (green pentagon in fig. 3, snapshot in fig. 2 d). For even higher levels of activity, the formation of chains is suppressed while global polarization persists, so that an oriented gas is observed (purple plus in fig. 3, snapshot in fig. 2 b).
When the system density exceeds the critical density of MIPS ($\Phi=0.57$, fig. 3 right), apolar ($\mu=0$) and weakly magnetic particles ($\mu<1$) also show ordered states like a network of chains (yellow square in fig. 3, snapshot in fig. 2 e) or a gas of chains (pink triangle). In case of apolar particles, increasing the activity results in cluster formation (brown triangle) which has previously been identified as MIPS [14] (and was confirmed here by a bimodal distributions of the local density). At high system density, network of chains (yellow square) and bands (red cross in fig. 3, snapshot in fig. 2 f) dominate the diagram of states.
III.2 Influence of an external field
@@
Figure 4: Collection of snapshots of the system at low density ($\Phi=0.13$) for intermediate self-propulsion speed $v_{0}=23$ for different magnetic field strengths $B$ (rows) and values for magnetic interaction strengths $\mu$ (columns). External magnetic field points from left to right. All snapshots have been taken at time $t=20$.
Figure 4 shows typical system configurations for three different values of the magnetic moments $\mu=$ 1.0 , 1.4 and 2.61.01.42.61.01.42.6start_ARG 1.0 end_ARG , start_ARG 1.4 end_ARG and start_ARG 2.6 end_ARG (columns from left to right) at these three field strengths ($B=$ 0.1 , 1.0 and 140.11.0140.11.014start_ARG 0.1 end_ARG , start_ARG 1.0 end_ARG and start_ARG 14 end_ARG in the top, middle, bottom row, respectively) for particles with moderate self-propulsion speed ($v_{0}=23$). The orientation of the particles is indicated by their color according to the color wheel. The magnetic field points from left to right, so alignment with the field can be seen by the cyan color of the particles. Systems with a weak external magnetic field (top row) display similar configurations to what was observed in systems without external magnetic field (compare with fig. 3). In systems with intermediate external magnetic fields (middle row), particles align with the direction of the external field (as seen by their cyan color). In addition, particles assemble into chains for strong external magnetic fields ($B=14$). These chains often have defects such as branches or small loops. Here, the average chain length is seen to increase with the magnetic moment of the particles, spanning the whole system size for strong magnetic interactions and strong external field strengths (for $\mu=2.6,\,B=14$, bottom right). These columnar ordered clusters are qualitatively similar to structures formed by non-equilibrium ferrofluids with external magnetic fields [23]. In contrast to systems without an external magnetic field, these columnar clusters mostly stay evenly separated and do not combine into connected networks of chains or into broad bands (at the low density simulated here, broad bands are seen for higher density, as discussed below). In fact, we observed in systems with strong magnetic interactions ($\mu=2.6$) that an increase of the external magnetic field strength causes broad bands to separate into thinner bands with fewer lanes or into individual chains of dipolar particles (compare the cases for $\mu=2.6$, $B=1.0$ and $\mu=2.6$, $B=14$ in fig. 4).
We analyzed the numerical simulations quantitatively by calculating the three global order parameters introduced above. They are plotted in fig. 5 for different $\mu$ and $v_{0}$ and for strong and weak magnetic field. Clustering (with $\left\langle\phi_{c}\right\rangle>0.5$) is largely absent at the low density used here and is only observed for the highest magnetic moment and a weak magnetic field (corresponding to the bands or networks of chains seen in the absence of a magnetic field). The strong magnetic field suppresses clustering also under these conditions (red lines).
@@
Figure 5: Three order parameters: fraction of largest cluster (top), polarization (center) and polymerization (bottom) against self-propulsion speed $v_{0}$ for four magnetic moments $\mu$ (colors). Solid color and solid lines show data for strong external magnetic field $B=14$, transparent colors and dashed lines show data with weak magnetic field strength $B=0.1$. System density was set to $\Phi=0.13$. Cut-off for cluster formation set at $\delta_{c}=1.28$.
@@
Figure 6: Diagrams of states for three different external magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.0140.11.014start_ARG 0.1 end_ARG , start_ARG 1.0 end_ARG and start_ARG 14 end_ARG (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at low density ($\Phi=0.13$).
@@
Using again the criteria based on the three order parameters, we classified the observed states of the system and summarized the results in a diagram of states, shown in fig. 6. The figure shows the diagrams of states for the three different magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.0140.11.014start_ARG 0.1 end_ARG , start_ARG 1.0 end_ARG and start_ARG 14 end_ARG (left to right) and for the density $\Phi=0.13$. Based on these diagrams we can make three key observations: First, disordered gas states (gray circles) are strongly suppressed with increasing magnetic field strength. In addition, the onset of the oriented gas state (purple plus) shifts to lower values of $\mu$. Second, intermediate as well as strong magnetic fields (center and right diagram) suppress the formation of networks of chains (yellow square) and instead promote the formation of oriented states, specifically the oriented gas (purple plus), oriented chains (green pentagon) and bands (red cross, for intermediate field strength). Third, and most importantly, strong external magnetic fields suppress the formation of bands. For strong external magnetic fields (right diagram), two states dominate the diagram, the oriented gas (purple plus) and oriented chains (green pentagon). The line separating these two states resembles the corresponding line separating the gas and chain state in small confined systems of dipolar active particles investigated in previous work [15].
@@
For strong magnetic fields, band patterns were largely absent (right panel in fig. 6). We only found bands for two parameter combinations ($\mu=2.6$, $v_{0}=200$ and $\mu$=4, $v_{0}=500$, indicated by the red crosses in fig. 6). In these cases we observed the following noteworthy behavior, shown by the snapshots in fig. 7. Dipolar active particles that have self-assembled into long chains transiently show oscillations. The amplitude of these oscillations was seen to grow towards their tail of a chain. We suspect that such oscillations might be caused by a buckling instability: In the chains, particles with strong dipolar interactions are packed rather closely. The combination of high swimming speeds and translational noise can then cause a temporary compression and stress along the chain that is resolved by buckling of the chain. When the dipolar interactions were reduced, long chains were seen to break into smaller segments. These breaks typically occurred near the tail of the chain, where the oscillations reach their maximum amplitude. This observation suggests that fragmentation induced by buckling may provide a mechanism for the coexistence of chains of different lengths. We note that the breaking of chains resembles the fragmentation of passive chains observed in magnetorheological suspensions driven by a rotating field [36, 37]. In our simulations these oscillating chains are sees as a transient state towards band formation. At some point, individual oscillating chains meet. Then, alignment with the neighboring particles suppresses oscillations and a band starts to form. Again, similar lateral aggregation of individual chains into bands and columnar lanes has been observed in magnetorheological fluids [22, 23, 24, 38].
@@
Figure 7: Series of snapshots showing oscillating chains for dipolar active particles with strong magnetic interactions $\mu=4$ and high self-propulsion speed $v_{0}=500$. Snapshots taken at times 26.46 , 26.56 and 26.6626.4626.5626.6626.4626.5626.66start_ARG 26.46 end_ARG , start_ARG 26.56 end_ARG and start_ARG 26.66 end_ARG (left to right), color wheel indicates orientations. External magnetic field points from left to right, magnetic field strength was set to $B=14$.
III.3 Effect of density
@@
To investigate the effect of the density of the system on the patterns formed by the active dipolar particles in a magnetic field, we simulated systems with two higher densities ($\Phi=0.23$ and $\Phi=0.57$) and repeated the calculation of the order parameters. The resulting state diagrams are shown in fig. 8 (for $\Phi=0.23$) and in fig. 9 (for $\Phi=0.57$).
@@
In systems with the intermediate density $\Phi=0.23$, we observed that intermediate external magnetic field strengths of $B=1$ do not suffice to suppress formation of complex networks (yellow squares, compare the middle panel in (fig. 8 with that for low density in fig. 6). Because of the increase in density and therefore a reduced mean particle distance, short chains may connect and form a complex network.
@@
Furthermore, the diagram of states indicates that strong external magnetic fields diminish the effect activity has in systems at intermediate density (fig. 8, right). Here, the line separating the oriented gas state (purple plus) and the oriented chains state (green pentagon) is independent of $v_{0}$. This observation is markedly different from the previous observations for systems without external magnetic fields and to systems at low density with external magnetic fields, where activity had a strong influence on the collective behavior and structure formation.
@@
Figure 8: Diagrams of states for three different external magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.0140.11.014start_ARG 0.1 end_ARG , start_ARG 1.0 end_ARG and start_ARG 14 end_ARG (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at intermediate density ($\Phi=0.23$). Cut-off for cluster formation set at $\delta_{c}=1.21$.
@@
For systems of high density (our case with $\Phi=0.57$), three phenomena compete with one another: polar order, chain formation and MIPS, as pointed out by Liao et al. [14]. The addition of an external magnetic field introduces global polar order and suppresses formation of complex networks of chains. Instead, we observed that strong external magnetic field strengths of $B=14$ clearly organize the system with high magnetic moments into columnar structures (see fig. 4 $\mu=2.6,B=14$). Because of the increase in density, these columnar structures are made up of multiple lanes in off-register alignment and not of single lanes, as seen in systems at low density.
@@
In the diagram of states, we observed that the increased density causes the previously suppressed band states (red crosses) to reappear. In fact, bands dominate the corresponding diagram of states for strong external magnetic fields and are independent of activity (fig. 9, right). Off-register band formation has been previously observed in passive dipolar systems with constant external magnetic fields [39, 40, 41]. In systems at high density, a clear discrimination between bands and columnar states can no longer be made. A limiting factor might be the finite system size, as observed in other systems of active particles [42].
@@
Figure 9: Diagrams of states for three different external magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.0140.11.014start_ARG 0.1 end_ARG , start_ARG 1.0 end_ARG and start_ARG 14 end_ARG (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at high density ($\Phi=0.57$). Cut-off for cluster formation set at $\delta_{c}=1.05$.
III.4 Columnar clusters